# The Human Labour of Data Work: Capturing Cultural Diversity through World Wide Dishes

### v2

@@ Acknowledgements
The authors would like to thank the following people for their feedback and insight during the development of World Wide Dishes: Kavengi Kitonga, Nari Johnson, Rida Qadri, Michael Leventhal, Rishi Vanukuru, Krishna Akhil Kumar Adavi, Julia Dean, Emma Ruttkamp-Bloem, Jordan Wirfs Brock and Morgan Scheuerman. We also acknowledge the significant contribution of those who contributed local expertise to the development of World Wide Dishes and the analyses we present here. We acknowledge these contributions in appendix A. This work has been supported by the Oxford Artificial Intelligence student society. SI was funded by the Ezoe Memorial Recruit Foundation until March 2024. TA is partially supported by ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. For computing resources, the authors are grateful for support from the Oxford Internet Institute, the OpenAI API Researcher Access Programme, as well as for the generous support from Jonathan Caton and the Google Cloud and Google’s Compute for Underrepresented Researchers Programme.
@@
Appendix A Contributors to the Dataset
@@ A.3. Contributors
Émilie Eliette-Caroline NGO Tjomb Assembe
@@ Appendix C Dataset Collection protocol depicted through Website Screenshots
Figure C.2. Condensed informed consent page. This page outlines important information that potential Contributors to World Wide Dishes must be aware of to make an informed decision about whether to contribute to the dataset. After reading through the provided list, they can choose to participate or not. The full informed consent is shown in Figure D.8 and Figure D.9. The condensed informed consent page was built after feedback from Contributors and Community Ambassadors that both the technical and essential meaning of the full form often got lost in translation to other languages.
Figure C.3. “About You” section, personal information form. After clicking on the “I’d like to participate and talk about food!” button on the informed consent page (Figure C.2), potential Contributors register or choose to continue as guests. Both options route the Contributors to the “About You” page, where they are asked to give information about their age and nationality.
@@
Figure C.5. About food section, continued. Contributors are given a series of questions to answer about the dish they are providing information about. The full list of data sought is shown in Table 1.
@@ E.1. Stakeholder statistics
Stakeholder Role Participant count National identities by continent Age range
Core Organisers The central organising team that coordinated the development and execution of World Wide Dishes. Core Organisers were also Community Ambassadors and Contributors 12 Africa, Asia, North America -
@@ E.3. Dish type tally
Table E.3. Number of dishes per type of meal. We ask Contributors to associate a dish with a specific type of meal. We accept that dishes can exist in different forms, so a one-to-many mapping is used. As such, dishes can be counted more than once.
@@ Appendix F Prompts used for Post-Mortem Reflections
The second author provided a series of prompts to the Core Organisers, who engaged in reflecting on the development and deployment of WWD. The second author sent these prompts in a series of text messages in a WhatsApp group chat with the COs.
@@ Appendix G Credits for Icons Used in Figures
  • ’Computer: Flaticon.com’. This figure has been designed using resources from Flaticon.com
  • ’Globe Icon: Flaticon.com’. This figure has been designed using resources from Flaticon.com
  • ’Database: Flaticon.com’. This figure has been designed using resources from Flaticon.com
  • ’Person Computer: Flaticon.com’. This figure has been designed using resources from Flaticon.com
  • ’Gears: Flaticon.com’. This figure has been designed using resources from Flaticon.com

### v3

@@ Acknowledgements
The authors would like to thank the following people for their feedback and insight during the development of World Wide Dishes: Kavengi Kitonga, Nari Johnson, Rida Qadri, Michael Leventhal, Rishi Vanukuru, Krishna Akhil Kumar Adavi, Julia Dean, Emma Ruttkamp-Bloem, Jordan Wirfs Brock and Morgan Scheuerman. We also acknowledge the significant contribution of those who contributed local expertise to the development of World Wide Dishes and the analyses we present here. We acknowledge these contributions in A. This work has been supported by the Oxford Artificial Intelligence student society. SI was funded by the Ezoe Memorial Recruit Foundation until March 2024. TA is partially supported by ELSA - European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. For computing resources, the authors are grateful for support from the Oxford Internet Institute, the OpenAI API Researcher Access Programme, as well as for the generous support from Jonathan Caton and the Google Cloud and Google’s Compute for Underrepresented Researcher’s Programme.
@@
Appendix A Contributors to the World Wide Dishes dataset
@@ A.3. Contributors
Émilie Eliette-Caroline sNGO Tjomb Assembe
@@ Appendix C Dataset collection protocol depicted through website screenshots
Figure C.2. Condensed informed consent page. This page outlines important information that potential Contributors to World Wide Dishes must be aware of to make an informed decision about whether to contribute to the dataset. After reading through the provided list, they can choose to participate or not. The full informed consent text is shown in Figure D.8 and Figure D.9. The condensed informed consent page was built after feedback from Contributors and Community Ambassadors that both the technical and essential meaning of the full form often got lost in translation to other languages.
Figure C.3. “About You” section. After clicking on the “I’d like to participate and talk about food!” button on the informed consent page (Figure C.2), potential Contributors register or choose to continue as guests. Both options route the Contributors to the “About You” page, where they are asked to give information about their age and nationality.
@@
Figure C.5. About food section, continued. Contributors are given a series of questions to answer about the dish for which they are providing information. The full list of data sought is shown in Table 1.
@@ E.1. Stakeholder statistics
Stakeholder Role Count National identities by continent Age range
Core Organisers The central organising team that coordinated the development and execution of World Wide Dishes. Core Organisers also acted as Community Ambassadors and Contributors 12 Africa, Asia, North America -
@@ E.3. Dish type tally
Table E.3. Number of dishes per type of meal. We asked Contributors to associate a dish with a specific type of meal. We accept that dishes can exist in different forms, so a one-to-many mapping is used. As such, dishes can be counted more than once.
@@ Appendix F Prompts used for post-mortem reflections
The second author provided a series of prompts to the Core Organisers, who reflected upon the development and deployment of WWD. The second author sent these prompts in a series of text messages in a WhatsApp group chat with the COs.
@@ Appendix G Credits for icons used in figures
  • ‘Computer: Flaticon.com’. This figure has been designed using resources from Flaticon.com..
  • ‘Globe Icon: Flaticon.com’. This figure has been designed using resources from Flaticon.com.
  • ‘Database: Flaticon.com’. This figure has been designed using resources from Flaticon.com.
  • ‘Person Computer: Flaticon.com’. This figure has been designed using resources from Flaticon.com.
  • ‘Gears: Flaticon.com’. This figure has been designed using resources from Flaticon.com.